# References.bib

### v1

@@ 2.2 Large Language Models
In recent years, large language models (LLMs) have grown in popularity as a driving force in artificial intelligence (AI), being used across various fields, such as trustworthiness [jin2024jailbreakzoo, deng2024deconstructing], scholarly document processing [zhuang2024understanding], signal processing [ruan2024s2e], quantum computing [liang2023unleashing], climate production [lin2023mmst, lin2023comprehensive], software engineering [zheng_survey_2024], and healthcare [he_large_2024] among multiple other learning environments. Zhao et al. [zhao_survey_2023] and Yang et al. [yang_harnessing_2023] define LLMs and pre-trained language models (PLMs) from the perspectives of model size and training approach. Generally speaking, PLMs refer to language models that are pre-trained on large amounts of general text data and then fine-tuned for specific tasks. LLMs are a kind of PLM. The key distinction is that LLMs are generally larger in scale with more parameters. These large language models have demonstrated the ability to learn universal representations of language, used in various natural language processing (NLP) tasks [huang_survey_2024], bolstering their applicability.
Figure 2: LLMs are a subset of AI. The focus of our paper is LLMs; however, we acknowledge AI due to the intertwined relationship between LLMs and AI.
LLMs are a sub-field of AI. Given this connection, discussions of blockchain for AI (BC4AI) research can also be applicable to LLMs. Figure 2 demonstrates the connections and following developments between AI, machine learning (ML), deep learning (DL), and LLMs. AI is largely discussed, yet we still see improvements in ML, such as FL that is better correlated to LLMs than general AI research. We then note DL as a subset of ML, showing advancements with neural network structures that are similar to how LLMs learn and operate. To define ML, we refer to Liu et al. [liu_when_2021] who state that ML is an automated learning process using algorithms and statistical models to efficiently perform specific tasks without the use of explicit instructions. DL is a sub-field of ML that utilizes multiple layers of artificial neurons to learn the latent space representation of the most basic form of data such as images, text, and speech signals. As described by Shafay et al. [shafay_blockchain_2023] deep learning has been able to achieve human-like accuracy, or even better accuracy than humans, on a variety of difficult tasks. Below, we elaborate upon the process of how models are trained and fine-tuned to become a usable, safe, and secure system.
@@ 2.2.1 Model Training
During the pre-training phase, the LLM is trained on a diverse, large dataset of textual data from various sources to learn the statistical properties of language. The LLM is equipped with a myriad of adjustable parameters, commonly reaching more than ten billion [huang_survey_2024]. Due to the huge model size and the vast amount of data used to train it, it is computationally challenging to successfully train a capable LLM, requiring distributed training algorithms for learning the model parameters [zhao_survey_2023]. Another crucial factor for LLM training is the data itself. Data that models are trained on come from a wide variety of sources, but the data itself may not be up to date [south_secure_2023]. To mitigate this shortcoming, recent advancements have introduced Retrieval-Augmented Generation (RAG), which is designed to augment and rectify the information returned by LLMs by consulting up-to-date online sources. The data that the LLM was trained on also has other deficiencies, like knowledge gaps in healthcare fields where data is private and restricted [ji_survey_2023]. Due to these knowledge gaps, the LLM may conjure up hallucinations where the model generates false information during prompting  [maynez_faithfulness_2020, openai_gpt-4_2024] because of a lack of relevant information. However, hallucinations may also occur with a plethora of data available as they are inherent problems in LLMs. Methods of preventing these hallucinations are elaborated on in  4.2.2. RAG can help rectify hallucinations, and fill in the gaps of data the LLM is missing, by using up-to-date and validated information from trustworthy online resources. This method of data retrieval introduces novel vulnerabilities since the information gathered by the retriever is largely un-audited and may contain poisoned data or data that can lead to unsafe responses from the LLMs.
@@ 2.2.2 Model Tuning and Utilization
After pre-training, the parameters of LLMs can be further updated by training on domain-specific datasets in downstream tasks. This process is known as fine-tuning (FT) [brown_language_2020]. A kind of fine-tuning method called supervised fine-tuning (SFT), aims to improve LLMs’ responsiveness to instructions, ensuring more desirable reactions involving three major components of instructions, inputs, and outputs. Inputs relate to prompting and the inputs depend on the instructions, similar to applications of open-ended generation in ChatGPT. By providing both inputs and outputs they form an instance, and multiple instances can exist for a single instruction [he_large_2024]. Among fine-tuning, there are a multitude of other training techniques within prompting the model such as instruction tuning and alignment tuning. By FT from a mixture of multi-task datasets formatted via natural language descriptions with the use of instruction tuning, LLMs are enabled to follow task instructions for new tasks without needing explicit examples, highlighting the ability of generalization for instruction following [zhao_survey_2023]. However, LLMs can demonstrate versatility, even without FT where they produce a phenomenon known as zero-shot learning, exhibiting the ability to perform tasks for which the model was never explicitly trained [brown_language_2020].
@@
Alignment tuning, a form of reinforcement learning, is used to promote the LLM to be a safe interactive machine. Since LLMs are trained to capture the data characteristics of uncurated pre-training corpora involving both high-quality and low-quality data, the LLM can generate toxic, biased, or harmful content for humans. To mitigate this problem, a FT process based on reinforcement learning from human feedback (RLHF) is used to align the LLM with safety values in order to make a trustworthy model [zhao_survey_2023]. The RLHF process ranks LLM outputs, with rewards scaled to positive and negative values. The LLM is then trained to produce highly-ranked responses and avoid low-ranked responses. In healthcare, RLHF provides advantages to the model such as improved accuracy and reliability through continuous feedback from medical professionals and customizes the interactions based on real clinical settings and patient needs [he_survey_2024]. These advanced training techniques improve LLM’s ability to generalize across tasks and improve their overall utility in various domains.
@@
Table 2: Differences in Definitions of Safety. There is no unifying definition of safety within the area of large language models. We see obvious agreement that models should be law-abiding, ethical, and non-violent in order to be safe, and as such these properties are strongly relevant to our definition of safety. However, beyond that point, there is generally a deviation between the authors’ respective definitions. This creates two further categories of terms, properties that are moderately relevant to safety and those that are weakly relevant. Questions of fairness, the informing ability of an LLM, and robustness are generally covered but not unanimously, and hence are moderately relevant, whereas privacy-preserving properties or non-sycophancy are rarely discussed in the current literature and are thus weakly relevant to safety. This dialogue between different modes of thought concerning what makes a large language model ”safe” heavily influences our definition of safety and our resulting discussion.
@@ 3 Research Methodology
First, and to allow a richer discussion centered around safety and security, we delineate between active and passive privacy within LLMs as introduced in Yan et al. [yan_protecting_2024].
@@
These definitions will serve a versatile role throughout this paper as building blocks for our contextualization of relevant and notable research efforts in the area of BC4LLMs. Moreover, they will also serve the community at large in helping to establish reliable and tangible properties of secure and safe large language models. They also will aid in establishing tighter definitions on finer-grained terms and ideas within BC4LLMs. For example, in Table 2 we provide definitions for terms found within our definition of safety to lessen the effect of the vague nature of some of the words. These definitions are backed by relevant examples found in the literature.
@@
Fairness LLMs ensuring a equitable environment for interaction, regardless of social identity. LLMs associating ”male” names with qualities of leadership, and ”female” names with qualities of amicability [wan_kelly_2023].
@@ 3.1 Research Approach and Limitations
Literature surveys often are limited in their depth and scope by unconscious factors that impact the authors’ ability to fairly select papers for review. To be transparent, and to aid researchers conducting similar or future reviews, we outline our research approach and its associated limitations. While conducting our research, we used the search engine Google Scholar, and several databases including ACM Computing Surveys, IEEE Xplore, SpringerLink, and arXiv. We chose these databases as they either produce quality research and contribute to the growth of interest in novel areas, or in the case of arXiv have the most up-to-date papers available. With Google Scholar, we used keyword searches such as ”blockchain for LLMs” and ”blockchain-based LLMs” as starting points for relevant, intriguing research papers. To solve problems concerning scope and interrelated domains of disparate areas, we gathered various applications of blockchain for AI, blockchain-enabled machine learning, federated learning, and deep learning tactics to apply them to LLMs. Lastly, of note is the fact that we were largely aided in this further research effort by a waterfall approach to finding research papers. That is, we found several foundational papers in the BC4LLM field, explored citations in those papers, and subsequently explored citations in those secondary papers. We continued investigating relevant citations in this waterfall fashion until we reasonably exhausted all relevant articles. Admittedly, this method of finding prominent research articles is limited in its natural tendency to develop blind spots to less well-known research articles or venues. However, in the spirit of a literature survey, we choose to focus on more established papers that more accurately capture the trends currently found in the space. For our exclusion criteria, we limited our research as follows: no duplicates; found articles from 2016 and above, excluding the original Merkle Tree paper [merkle_digital_1988]; no Masters or Ph.D. theses; and only studies written in English.
@@ 3.2 Model of Threat Categorization
There exist a wide variety of threats which affect LLMs. Oftentimes, many of these threats originate from the nature of LLMs acting as AI systems. In Figure 3 we refer to these vulnerabilities similarly to Yao et al [yao_survey_2024] who categorized the most extensively discussed LLM vulnerabilities and AI-inherent vulnerabilities together, yet also included external threats under non-AI inherent vulnerabilities. We contribute further by applying these vulnerabilities to each respective process within developing a large language model and tie these to respective applications of blockchain. Beginning with the training process, LLMs are prone to threats such as data poisoning and backdoor attacks. As defined in Yao et al. [yao_survey_2024] data poisoning is where attackers influence the training process by injecting malicious data into the training set, introducing vulnerabilities within the security and effectiveness of the model. Following the trend of poisoned data, there can be backdoor attacks implemented on the training data, as defined in Li et al. [li_blockchain-based_2023] who categorize backdoor attacks into attacks on training data and attacks on local models. The backdoor attacks on training data are further divided into attacks based on label flipping and attacks based on planting triggers. Attacks based on label flipping focus on manipulating the labels, whereas attacks on planting triggers modify the input data and labels, effectively constructing an adversarial sample. Then, attacks on local models are further divided into attacks based on modifications to the training process and attacks based on manipulating the trained model [li_blockchain-based_2023]. The backdoor attacks can be applicable to both the training and prompting phases of LLMs when using this distinction.
@@
RAG attacks have a variety of issues, including privacy issues [zeng_good_2024][anderson_is_2024] and knowledge poisoning attacks [zou_poisonedrag_2024]. For RAG specifically, Xue et al. [xue_badrag_2024] propose BadRAG to identify security vulnerabilities, exposing direct attacks on the retrieval phase from semantic triggers, and uncovering indirect attacks on the generative phase of LLMs that were caused by a contaminated corpus. These RAG-specific attacks and defenses are elaborated on in section 4.1.2. When interacting with a LLM, we see AI’s inherent vulnerabilities become relevant, as in Yao et al. [yao_survey_2024] note that LLMs are fundamentally AI models themselves. We focus on the prevalent adversarial attacks that malicious users may use to tamper with the LLM, attempts to find out sensitive information, or try to ruin the system entirely. We recognize jailbreaking and prompt injection as two separate but similar types of adversarial attacks that are initiated within prompting. For instance, jailbreaking prompts are designed to bypass the restrictions set by service providers during model alignment or other containment approaches [shayegani_survey_2023]. Whereas prompt injection attacks aim to override an LLM’s original prompt and directs it to follow a set of malicious instructions, leading to erroneous advice or unauthorized data leakage [liu_prompt_2024]. In our Sections 4.2.1, 4.2.2 we discuss instances of misinformation and passive privacy leakage addressed as safety concerns. Note that from above, we included backdoor attacks based on modifications to the trained model in prompting, since these backdoor attacks can still happen after the model has been trained [li_blockchain-based_2023].
@@
Other attacks relating to LLMs and AI are membership inference attacks (MIA), relating to a type of privacy attack where some malicious user, given access to the model, is able to determine whether a given point was used to train that model with high accuracy [neel_privacy_2024]. However, Neel and Chang [neel_privacy_2024] state this attack is more related to information about the training point data leaking through the model, and that malicious users must have access to a candidate point in order to run the attack. Therefore this attack is more prevalent with passive privacy, highlighting the need for preventing data leakage. Similar attacks are user inference attacks that seek to gain knowledge or insights about the model or data’s characteristics, often by observing the model’s responses or behavior [yao_survey_2024].
@@
Lastly, we explore denial of service (DoS) attacks and supply chain vulnerabilities. From Yao et al. [yao_survey_2024] a DoS attack is a type of cyber attack that aims to exhaust computational resources, resulting in latency or making the technology resources unavailable. We focus on distributed denial of service attacks (DDoS) which is a type of DoS attack where requests flood the system, attacking simultaneously from multiple sources on the network [elmamy_survey_2020]. Yao et al. [yao_survey_2024] also defined LLM supply chain vulnerabilities, referring to the risks in the lifecycle of LLM applications that may occur from using vulnerable components or services, including third party plugins that may be used to steal chat histories, access private information, and or execute code on a user’s machine [yao_survey_2024]. All of these security vulnerabilities are substantial threats to LLMs that need to be mitigated or prevented. Possible methods of defense are discussed within Section 4.1, using current blockchain frameworks and experiments for these security problems, as listed by each developmental phase of the LLM, AI inherent threats, and supply chain issues.

### v2

@@ 2.2 Large Language Models
In recent years, large language models (LLMs) have emerged as a pivotal force in artificial intelligence (AI), contributing to widespread applications across diverse fields, such as trustworthiness [deng2024deconstructing, jiang2024robustkv, jin2024jailbreakzoo, liu2024buckle], scholarly document processing [zhuang2024understanding], signal processing [ruan2024s2e], quantum computing [liang2023unleashing], climate production [lin2023comprehensive, lin2023mmst], software engineering [zheng_survey_2023], and healthcare [he_large_2024] among multiple other learning environments. Zhao et al. [zhao_survey_2023] and Yang et al. [yang_harnessing_2023] define LLMs and pre-trained language models (PLMs) from the perspectives of model size and training approach. Generally speaking, PLMs refer to language models that are pre-trained on large amounts of general text data and then fine-tuned for specific tasks. LLMs are a kind of PLM. The key distinction is that LLMs are generally larger in scale with more parameters. These large language models have demonstrated the ability to learn universal representations of language, used in various natural language processing (NLP) tasks [huang_survey_2024], bolstering their applicability.
Figure 2: The connection among AI, Machine Learning (ML), Deep Learning (DL), and LLMs.
We discuss connections and following developments between AI, machine learning (ML), deep learning (DL), and LLMs in Figure 2. AI refers to a broad technique that aims at simulating human intelligence, encompassing a variety of approaches and methods. Machine Learning (ML) is a subfield of AI that develops algorithms and statistical models to automatically learn from data and efficiently perform specific tasks without the use of explicit instructions [liu_when_2021]. Deep Learning (DL) is a subset of ML that utilizes multi-layered neural networks to learn latent representation on various tasks [shafay_blockchain_2023]. LLMs are one of the popular applications using cutting-edge DL models, advancing natural language understanding and generation at the human level. In the following subsections, we elaborate on the process of how LLMs are pre-trained and utilized as safe and powerful AI systems.
@@ 2.2.1 Model Training
During the pre-training phase, the LLM is trained on a diverse, large dataset of textual data from various sources to learn the statistical properties of language. The LLM is equipped with a myriad of adjustable parameters, commonly reaching more than ten billion [huang_survey_2024]. Due to the huge model size and the vast amount of data used to train it, it is computationally challenging to successfully train a capable LLM, requiring distributed training algorithms for learning the model parameters [zhao_survey_2023]. Another crucial factor for LLM training is the data itself. Data that models are trained on come from a wide variety of sources, but the data itself may not be up to date [south_secure_2023]. To mitigate this shortcoming, recent advancements have introduced Retrieval Augmented Generation (RAG), which is designed to augment and rectify the information returned by LLMs by consulting up-to-date online sources. The data that the LLM was trained on also has other deficiencies, like knowledge gaps in healthcare fields where data is private and restricted [ji_survey_2023]. Due to these knowledge gaps, the LLM may conjure up hallucinations where the model generates false information during prompting [maynez_faithfulness_2020, openai_gpt-4_2024] because of a lack of relevant information. However, hallucinations may also occur with a plethora of data available as they are inherent problems in LLMs. Methods of preventing these hallucinations are elaborated in Section 4.2.2. RAG can help rectify hallucinations, and fill in the gaps of data the LLM is missing, by using up-to-date and validated information from trustworthy online resources. This data retrieval method introduces novel vulnerabilities since the information gathered by the retriever is largely unaudited and may contain poisoned data or data that can lead to unsafe responses from the LLMs.
@@ 2.2.2 Model Tuning and Utilization
After pre-training, the parameters of LLMs can be further updated by training on domain-specific datasets in downstream tasks. This process is known as fine-tuning (FT) [brown_language_2020]. A kind of fine-tuning method called supervised fine-tuning (SFT), aims to improve LLMs’ responsiveness to instructions, ensuring more desirable reactions involving three major components of instructions, inputs, and outputs. Inputs relate to prompting and the inputs depend on the instructions, similar to applications of open-ended generation in ChatGPT. By providing both inputs and outputs they form an instance, and multiple instances can exist for a single instruction [he_large_2024]. Among fine-tuning, other training techniques within model prompting include instruction tuning and alignment tuning. By FT from a mixture of multi-task datasets formatted via natural language descriptions with the use of instruction tuning, LLMs are enabled to follow task instructions for new tasks without needing explicit examples, highlighting the ability of generalization for instruction following [zhao_survey_2023]. However, LLMs can demonstrate versatility, even without FT where they produce a phenomenon known as zero-shot learning, exhibiting the ability to perform tasks for which the model was never explicitly trained [brown_language_2020].
@@
Alignment tuning, equipped with reinforcement learning, is used to enhance LLMs to be safe interactive models. Since LLMs are trained to capture the data characteristics of uncurated pre-training corpora involving both high-quality and low-quality data, the LLM can generate toxic, biased, or harmful content for humans. To mitigate this problem, an FT process based on reinforcement learning from human feedback (RLHF) is used to align the LLM with the outcomes that satisfy human values [zhao_survey_2023]. The RLHF process ranks LLM outputs, with rewards scaled to positive and negative values. The LLM is then trained to produce highly-ranked responses and avoid low-ranked responses. In healthcare, RLHF provides advantages to the model such as improved accuracy and reliability through continuous feedback from medical professionals, and customizes the interactions based on real clinical settings and patient needs [he_survey_2024]. These advanced training techniques improve LLM’s ability to generalize across tasks and improve their overall utility in various domains.
@@
Table 2: Differences in Definitions of Safety. There is no unifying definition of safety within the area of large language models. We see obvious agreement that models should be law-abiding, ethical, and non-violent in order to be safe, and as such these properties are strongly relevant to our definition of safety. However, beyond that point, there is generally a deviation between the authors’ respective definitions. This creates two further categories of terms, properties that are moderately relevant to safety and those that are weakly relevant. Questions of fairness, the informing ability of an LLM, and robustness are generally covered but not unanimously, and hence are moderately relevant, whereas privacy-preserving properties or non-sycophancy are rarely discussed in the current literature and are thus weakly relevant to safety. This dialogue between different modes of thought concerning what makes a large language model “safe” heavily influences our definition of safety and our resulting discussion.
@@ 3 Research Methodology
First, and to allow a richer discussion centered around safety and security, we delineate between active and passive privacy within LLMs as introduced in [yan_protecting_2024].
@@
These definitions will serve a versatile role throughout this paper as building blocks for our contextualization of relevant and notable research efforts in BC4LLMs. Besides, they will serve the community at large in helping to establish reliable and tangible properties of secure and safe large language models. Furthermore, They will help establish tighter definitions of finer-grained terms and ideas within BC4LLMs. For example, in Table 2, we provide definitions for terms found within our definition of safety to lessen the effect of the vague nature of some of the words. These definitions are backed by relevant examples found in the literature.
@@
Fairness LLMs ensuring a equitable environment for interaction, regardless of social identity. LLMs associating “male” names with qualities of leadership, and “female” names with qualities of amicability [wan_kelly_2023].
@@ 3.1 Research Approach and Limitations
Literature surveys often are limited in their depth and scope by unconscious factors that impact the authors’ ability to fairly select papers for review. To be transparent, and to aid researchers conducting similar or future reviews, we outline our research approach and its associated limitations. While conducting our research, we used the search engine, Google Scholar, and several databases, including ACM Computing Surveys, IEEE Xplore, SpringerLink, and arXiv. We chose these databases as they either produce quality research and contribute to the growth of interest in novel areas, or in the case of arXiv have the most up-to-date papers available. With Google Scholar, we used keyword searches such as “blockchain for LLMs” and “blockchain-based LLMs” as starting points for relevant, intriguing research papers. To solve problems concerning the scope and interrelated domains of disparate areas, we gathered various applications of blockchain for AI, blockchain-enabled machine learning, federated learning, and deep learning tactics to apply them to LLMs. Lastly, of note is the fact that we were largely aided in this further research effort by a waterfall approach to finding research papers. That is, we found several foundational papers in the BC4LLM field, explored citations in those papers, and subsequently explored citations in those secondary papers. We continued investigating relevant citations in this waterfall fashion until we reasonably exhausted all relevant articles. Admittedly, this method of finding prominent research articles is limited in its natural tendency to develop blind spots to less well-known research articles or venues. However, in the spirit of a literature survey, we choose to focus on more established papers that more accurately capture the trends currently found in the space. For our exclusion criteria, we limited our research as follows: no duplicates; found articles from 2016 and above, excluding the original Merkle Tree paper [merkle_digital_1988]; no Masters or Ph.D. theses; and only studies written in English.
@@ 3.2 Model of Threat Categorization
There exists a wide variety of threats that affect LLMs. Oftentimes, many of these threats originate from the nature of LLMs acting as AI systems. In Figure 3, we refer to these vulnerabilities similarly to that discussed in [yao_survey_2024], which categorized the most LLM vulnerabilities and AI-inherent vulnerabilities together, yet also included external threats under non-AI inherent vulnerabilities. We contribute further by applying these vulnerabilities to each respective process within developing a large language model and tying these to respective applications of blockchain. Beginning with the training process, LLMs are prone to threats such as data poisoning and backdoor attacks. As defined by Yao et al. [yao_survey_2024], data poisoning is where attackers influence the training process by injecting malicious data into the training set, introducing vulnerabilities within the security and effectiveness of the model. Following the trend of poisoned data, there can be backdoor attacks implemented on the training data, as defined by Li et al. [li_blockchain-based_2023], who categorize backdoor attacks into attacks on training data and attacks on local models. The backdoor attacks on training data are further divided into attacks based on label flipping and attacks based on planting triggers. Attacks based on label flipping focus on manipulating the labels, whereas attacks on planting triggers modify the input data and labels, effectively constructing an adversarial sample. Then, attacks on local models are further divided into attacks based on modifications to the training process and attacks based on manipulating the trained model [li_blockchain-based_2023]. The backdoor attacks can be applicable to both the training and prompting phases of LLMs when using this distinction.
@@
RAG attacks have a variety of issues, including privacy issues [zeng_good_2024, anderson_is_2024] and knowledge poisoning attacks [zou_poisonedrag_2024]. For RAG specifically, Xue et al. [xue_badrag_2024] propose BadRAG to identify security vulnerabilities, exposing direct attacks on the retrieval phase from semantic triggers, and uncovering indirect attacks on the generative phase of LLMs that were caused by a contaminated corpus. These RAG-specific attacks and defenses are elaborated on in Section 4.1.2. When interacting with an LLM, AI’s inherent vulnerabilities become evident, as highlighted in [yao_survey_2024], since LLMs are fundamentally AI models themselves. We focus on the prevalent adversarial attacks that malicious users may use to tamper with the LLM, attempt to find out sensitive information, or ruin the system entirely. We recognize jailbreaking and prompt injection as two separate but similar types of adversarial attacks that are initiated within prompting. For instance, jailbreaking prompts are designed to bypass the restrictions set by service providers during model alignment or other containment approaches [shayegani_survey_2023]. Prompt injections aim to override an LLM’s original prompt and direct it to follow a set of malicious instructions, leading to erroneous advice or unauthorized data leakage [liu_prompt_2024]. In Sections 4.2.1 and 4.2.2, we discuss instances of misinformation and passive privacy leakage addressed as safety concerns. Note that we include backdoor attacks based on modifications to the trained model in prompting since these backdoor attacks can still happen after model training [li_blockchain-based_2023].
@@
Another relevant attack is a membership inference attack (MIA), a type of privacy attack where some malicious users, given access to the model, can determine whether a given point was used to train that model with high accuracy [neel_privacy_2024]. However, Neel and Chang [neel_privacy_2024] state that this attack is more related to information about the training point data leaking through the model, and that malicious users must have access to a candidate point in order to run the attack. Therefore, this attack is more prevalent with passive privacy, highlighting the need to prevent data leakage. Similar attacks are user inference attacks that seek to gain knowledge or insights about the model or data’s characteristics, often by observing the model’s responses or behavior [yao_survey_2024].
@@
Last but not least, we explore denial of service (DoS) attacks and supply chain vulnerabilities. Yao et al. [yao_survey_2024] describe DoS attacks as a type of cyber attack that aims to exhaust computational resources, resulting in latency or making the technology resources unavailable. In this survey, we focus on distributed denial of service attacks (DDoS), a type of DoS attack where requests flood the system, attacking simultaneously from multiple sources on the network [elmamy_survey_2020]. Yao et al. [yao_survey_2024] also defined LLM supply chain vulnerabilities as the risks in the lifecycle of LLM applications that may occur from using vulnerable components or services, including third-party plugins that may be used to steal chat histories, access private information, and or execute code on a user’s machine. All of these security vulnerabilities are substantial threats to LLMs that need to be mitigated or prevented. Possible methods of defense are discussed in Section 4.1, using current blockchain frameworks and experiments for these security problems, as listed by each developmental phase of the LLM, AI inherent threats, and supply chain issues.